OpenStudy (anonymous):

Identify the vertex, focus, and directrix of the parabola with the equation x^2 – 6x – 8y + 49 = 0

OpenStudy (anonymous):

@TuringTest can you help me please??

OpenStudy (accessdenied):

Have you tried putting this equation into the vertex form?

OpenStudy (anonymous):

how do you do that?

OpenStudy (anonymous):

@AccessDenied ?

OpenStudy (accessdenied):

You know the vertex form looks like this? \( y = a(x - h)^2 + k \) just to make sure you are familiar with what I am referring to.

OpenStudy (anonymous):

yes

OpenStudy (accessdenied):

So, we could complete the square here: \( \color{Green}{x^2 - 6x} - 8y + 49 = 0 \) Which will give us that \( (x - h)^2\) part. Then rearrange so that the y term is alone on the other side of the equation.

OpenStudy (anonymous):

ok..then what?

OpenStudy (accessdenied):

If you do that part, you will obtain the vertex form: \(y = a (x- h)^2 + k\) From this form, the vertex is directly given by (h, k). And a = 1/(4c), where c is the distance to either focus or directrix.

OpenStudy (accessdenied):

You use the vertex as the basis point. The distance from vertex to focus is that value of c, and we are moving a vertical distance because the parabola is opening upwards as the focus lies on the axis of symmetry (vertical line through vertex).  Does that make sense? Do I need to cover anything more in-depth? :)